0.229 g sample of carbon dioxide, CO2, has a volume of 539 mL and a pressure of 458 mmHg .

1.What is the temperature of the gas in kelvins?

2.What is the temperature of the gas in degrees Celsius?
1
Expert's answer
2018-04-11T08:11:34-0400
Answer:- Using ideal gas equation
PV=nRT=mRT/M
T=PVM/mR
P=458/760=0.6026atm
V=539/1000=0.539 l
T=0.6026*0.539*44/0.8321*0.229=75.01k

In celcius
T=75.01-273.1
T=-198.14 degree celcius